# Enhancing Physics-Informed Neural Networks Through Feature Engineering

### v2

@@ 7 Discussion and Future Work
While SAFE-NET demonstrates significant improvements in conditioning and convergence for smooth PDEs, several avenues remain to improve feature engineering in PINNs. The first is Multi-stage feature learning, following (Wang & Lai, 2023), a multi-stage approach that further boosts performance by iteratively learning a best fit on the residual error, using SAFE-NET as a base learner could be developed. Another direction is better incorporating physical priors. Enforcing physical laws such as conservation principles and symmetries into the feature engineering process could improve the performance of SAFE-NET on complex problems. Features based on concepts like Noether’s theorem or Hamiltonian mechanics could provide stronger inductive biases. Finally, it would be interesting to also consider non-Fourier features like radial basis functions for handling sharp gradients, which could allow SAFE-NET to perform well on a wider range of problems. The challenge here lies in maintaining numerical stability when combining these different types of features.
@@ Appendix A Additional Experimental Setups
For all experiments, we maintain consistent architectural and training configurations across all methods to ensure fair comparison. Unless stated otherwise in the method-specific sections below, all PINN-based models utilize a fully connected neural network architecture with 4 hidden layers, each containing 50 neurons, and employ the $\tanh$ activation function. Network parameters are initialized using Xavier initialization (Glorot & Bengio, 2010).
@@ A.1.2 Feature Generator
$\phi_{1}^{(i)}(x,t)$ $=\text{coeff}_{1}^{(i)}\cdot\sin(\omega_{x}^{(i)}\cdot x)\cdot\cos(\omega_{t}^{(i)}\cdot t),$ (17)
$\phi_{2}^{(i)}(x,t)$ $=\text{coeff}_{2}^{(i)}\cdot\sin(\omega_{t}^{(i)}\cdot t)\cdot\cos(\omega_{x}^{(i)}\cdot x),$ (18)
$\phi_{3}^{(i)}(x,t)$ $=\text{coeff}_{3}^{(i)}\cdot\sin(\omega_{x}^{(i)}\cdot x)\cdot\sin(\omega_{t}^{(i)}\cdot t),$ (19)
$\phi_{4}^{(i)}(x,t)$ $=\text{coeff}_{4}^{(i)}\cdot\cos(\omega_{x}^{(i)}\cdot x)\cdot\cos(\omega_{t}^{(i)}\cdot t).$ (20)
Here, $\omega_{x}^{(i)}$ and $\omega_{t}^{(i)}$ are trainable frequencies, and $\text{coeff}_{j}^{(i)}$ are trainable coefficients for $j=1,2,3,4$ and $i=1,\ldots,N$ where $N$ is the number of Fourier feature sets.
@@
  • Heat equation: Initial condition $u(x,y,0)=\sin(20\pi x)\sin(\pi y)$ and homogeneous boundary condition $u(x,y,t)=0$ on $\partial\Omega$ for the unit square domain $[0,1]^{2}$ suggest features that satisfy the boundary conditions while capturing the initial structure: $\{\sin(20\pi x),\sin(\pi y)\}$ and their products $\{\sin(20\pi x)\sin(\pi y)\}$
  • Burgers equation: Initial condition $u(x,0)=-\sin(\pi x)$ with homogeneous Dirichlet boundaries $u(-1,t)=u(1,t)=0$ on domain $[-1,1]$ provides the feature $\{-\sin(\pi x)\}$.
@@
  • Allen-Cahn equation: Initial condition $u(0,x)=x^{2}\cos(\pi x)$ with periodic boundary conditions $u(t,x-1)=u(t,x+1)$ on domain $[-1,1]$ suggests features based on the polynomial-trigonometric structure: $\{x^{2},\cos(\pi x),x^{2}\cos(\pi x)\}$.
  • Navier-Stokes equation: For the back-step flow geometry $[0,4]\times[0,2]\setminus([0,2]\times[1,2])$ with no-slip boundary conditions and inlet condition $u_{x}=4y(1-y)$, features are derived from the inlet profile and domain constraints: $\{y(1-y),y,1-y\}$ and polynomial combinations that respect the geometric constraints.
@@ A.1.4 Parameter Initialization
  • Fourier Frequencies: Initialized as $\omega_{x}^{(i)}=\omega_{t}^{(i)}=i\pi$ for $i=1,2,\ldots,N$, where $N$ is the number of Fourier feature sets. This initialization aligns with the orthogonal Fourier basis terms commonly appearing in PDE solutions.
  • Fourier Coefficients: All coefficients $\text{coeff}_{j}^{(i)}$ are initialized to 1, providing equal weight to all Fourier modes initially.
@@ A.2.1 PINN and FLS-PINN
PINN: We implement the standard Physics-Informed Neural Network as proposed by Raissi et al. (Raissi et al., 2019b). This serves as our primary baseline, utilizing the vanilla MLP architecture with coordinate inputs directly fed to the network without any feature engineering or positional encoding. The method relies solely on the network’s inherent ability to learn complex mappings between spatial-temporal coordinates and the solution field through the physics-informed loss function. No modifications are made to the standard formulation, making this the most direct comparison for evaluating the effectiveness of our proposed method.
FLS-PINN: The First Layer Sine PINN represents a modification inspired by the SIREN architecture (Sitzmann et al., 2020), where we replace the $\tanh$ activation function in the first hidden layer with the $\sin$ activation function, while maintaining $\tanh$ activations in all subsequent layers. This hybrid approach aims to leverage the beneficial properties of sinusoidal activations for coordinate-based inputs, particularly their ability to represent high-frequency components, while preserving the stability characteristics of $\tanh$ activations in deeper layers. The first layer weights are initialized using a uniform distribution $\mathcal{U}(-1/d_{\text{in}},1/d_{\text{in}})$ where $d_{\text{in}}$ is the input dimension, following the SIREN initialization scheme, while subsequent layers maintain Xavier initialization. This configuration has been shown to improve convergence behavior for certain classes of PDEs, particularly those involving high-frequency solutions or sharp gradients.
@@ A.2.2 W-PINN
From (Wang et al., 2022a), W-PINN (Weighted Physics-Informed Neural Network) is a variant of PINN that addresses spectral bias and imbalanced convergence rates in multi-term loss functions through adaptive weight calibration. For a PDE defined as:
@@
We refer the reader to (Wang et al., 2021b) for details of these calculations and definitions. The key insight of W-PINN is that the eigenvalues of the NTK matrices $\mathbf{K}_{uu}$ and $\mathbf{K}_{rr}$ characterize the convergence rates of the boundary and residual losses respectively. The method proposes adapting the weights according to:
@@
Implementation Details. In our experiments, we run W-PINN following the setup from (Wang et al., 2022b). The method down-weights the residual loss term to reduce its dominance and improve the fitting of boundary conditions. Specifically, we tested different weighting configurations for the residual loss term from the set $\{10^{-1},10^{-3},10^{-4}\}$ while maintaining the boundary condition weight at the standard value. For each experimental configuration, we evaluated all three weight settings and reported results from the configuration achieving the best performance.
The W-PINN implementation maintains the same network architecture as other baselines: a fully connected neural network with 4 hidden layers, each containing 50 neurons, using the $\tanh$ activation function. Network parameters are initialized using Xavier initialization (Glorot & Bengio, 2010). The adaptive weight computation is performed at regular intervals during training to update the loss function weights based on the current Neural Tangent Kernel eigenvalue analysis.
@@ A.2.3 RBA-PINN
For the RBA-PINN baseline (Anagnostopoulos et al., 2023), we implement the residual-based attention scheme coupled with Fourier feature embeddings (RBA+Fourier configuration) to leverage the method’s demonstrated strengths while maintaining computational efficiency. Based on the original paper’s ablation studies, this configuration represents the most effective combination, where RBA alone achieves $3.16\times 10^{-3}$ relative $L^{2}$ error on the Allen-Cahn equation, while the RBA+Fourier variant improves performance by two orders of magnitude.
Network Architecture. Following the original implementation (Anagnostopoulos et al., 2023), we employ a fully connected neural network with 6 hidden layers and 50 neurons per layer, utilizing the $\tanh$ activation function. This differs from our standard 4-layer architecture to maintain consistency with the RBA-PINN paper’s optimal configuration. Network weights are initialized using Xavier initialization (Glorot & Bengio, 2010).
@@
Fourier Feature Embeddings. To enforce periodic boundary conditions as hard constraints, we implement one-dimensional Fourier feature embeddings following (Anagnostopoulos et al., 2023). For problems with periodic boundary conditions in the spatial domain, the input $x$ is augmented with Fourier features:
@@
[…]
@@ A.2.4 RFF
The Random Fourier Feature PINN (RFF) builds upon the theoretical foundations of Bochner’s theorem and Neural Tangent Kernel (NTK) theory by employing random Fourier feature mappings as coordinate embeddings before the input layer of the neural network (Tancik et al., 2020a; Wang et al., 2021b). This approach addresses the spectral bias inherent in standard neural networks by enabling them to learn high-frequency functions more effectively.
@@
Implementation Details. Following the methodology established in (Wang et al., 2021b), we implement RFF using 128 features ($m=64$, resulting in $1282m=$ 128 total features after cosine and sine transformations). The frequency scaling parameter $\sigma$ is chosen based on the expected frequency characteristics of the solution domain. Specifically, we adopt a coordinate-dependent scaling strategy:
@@
The random matrix $B$ is sampled once during initialization and remains fixed throughout training, ensuring deterministic feature mappings. We utilize the random-fourier-features-pytorch library implementation (Long, 2021) for computational efficiency.
@@
Hyperparameter Selection. The frequency parameter $\sigma$ requires careful tuning based on problem characteristics. Following the analysis in (Wang et al., 2021b), we conducted preliminary experiments with $\sigma\in\{1,10,50,100,200\}$ for spatial coordinates and $\sigma\in\{1,5,10,20\}$ for temporal coordinates. The selected values ($\sigma=200$ for space, $\sigma=10$ for time) consistently provided the best performance across our benchmark problems.
@@ A.2.5 RBF and RBF-P
RBF (Radial Basis Functions PINN): We implement the RBF-based feature mapping method from Zeng et al. (Zeng et al., 2024b), which introduces non-Fourier positional embedding as an alternative to commonly used Fourier-based feature mappings. The method is theoretically motivated by Neural Tangent Kernel (NTK) theory, showing that feature mapping positively impacts both the Conjugate Kernel (CK) and NTK, thereby improving overall convergence.
@@
$\Phi(x)=\frac{w_{i}\phi(|x-c_{i}|)}{\sum_{j=1}^{m}w_{j}\phi(|x-c_{j}|)}$ (36)
where $x\in\mathbb{R}^{n}$ is the input coordinate, $c_{i}\in\mathbb{R}^{m}$ are the RBF centers, and $w_{i}$ are the trainable weights. We employ Gaussian RBFs as the radial basis function:
$\phi(r)=e^{-\frac{r^{2}}{\sigma^{2}}}$ (37)
where $r=|x-c_{i}|$ represents the Euclidean distance between input and RBF center, and $\sigma$ controls the bandwidth.
@@
[…]
@@ A.3 Optimizers
As discussed in the main text, we employ a systematic approach to optimizer selection, motivated by recent loss-landscape studies showing that first-order methods like Adam excel at escaping saddle points but struggle in ill-conditioned valleys, while quasi-Newton methods like L-BFGS provide superior local conditioning but are prone to stalling at saddles (Rathore et al., 2024). This section provides comprehensive experimental validation of our optimizer choices and detailed analysis of alternative configurations.
@@ A.3.1 Experimental Configurations
We conducted extensive experiments across multiple optimizer combinations to validate our scheduling choices. All experiments with hybrid Adam and L-BFGS combination optimizaers maintain a total training budget of 40k iterations across the tested PDE benchmarks. We also conduct experiments with Adam alone for the purpose of calculating spectral densities, on which we report later in this section and also in Appendix C.2 and C.3. We list the configurations already explained in the main text below and move on to studying other possible configurations and their results.
@@ A.3.2 Experimental Results
[…]

### v3

@@ 7 Discussion and Future Work
Fourier features align best with low-to-moderate-dimensional PDEs demonstrating smoother behavior; for strongly nonlinear or shock-forming problems, non-Fourier features (e.g., RBF-P) are often more suitable, consistent with our results. While SAFE-NET demonstrates significant improvements in conditioning and convergence for PDEs whose solutions exhibit predominantly smooth behavior (including piecewise-smooth profiles without strong shocks), several avenues remain to improve feature engineering in PINNs. The first is multi-stage feature learning, following wang2023multistageneuralnetworksfunction, a multi-stage approach that further boosts performance by iteratively learning a best fit on the residual error, using SAFE-NET as a base learner could be developed. Another direction is better incorporating physical priors. Enforcing physical laws such as conservation principles and symmetries into the feature engineering process could improve the performance of SAFE-NET on complex problems. Features based on concepts like Noether’s theorem or Hamiltonian mechanics could provide stronger inductive biases. Finally, it would be interesting to also consider non-Fourier features like radial basis functions for handling sharp gradients, which could allow SAFE-NET to perform well on a wider range of problems. The challenge here lies in maintaining numerical stability when combining these different types of features. Most importantly, extending feature engineering to truly high‑dimensional problems is important future work, but it presents very different challenges that are currently beyond the reach of PINN-style methods.
@@ Appendix A Additional Experimental Setups
For all experiments, we maintain consistent architectural and training configurations across all methods to ensure fair comparison. Unless stated otherwise in the method-specific sections below, all PINN-based models utilize a fully connected neural network architecture with 4 hidden layers, each containing 50 neurons, and employ the $\tanh$ activation function. Network parameters are initialized using Xavier initialization glorot2010understanding.
@@ A.1.2 Feature Generator
$\phi_{1}^{(\ell)}(x,t)$ $=\text{coeff}_{1}^{(\ell)}\cdot\sin(\omega_{x}^{(\ell)}\cdot x)\cdot\cos(\lambda_{t}^{(\ell)}\cdot t),$ (17)
$\phi_{2}^{(\ell)}(x,t)$ $=\text{coeff}_{2}^{(\ell)}\cdot\sin(\lambda_{t}^{(\ell)}\cdot t)\cdot\cos(\omega_{x}^{(\ell)}\cdot x),$ (18)
$\phi_{3}^{(\ell)}(x,t)$ $=\text{coeff}_{3}^{(\ell)}\cdot\sin(\omega_{x}^{(\ell)}\cdot x)\cdot\sin(\lambda_{t}^{(\ell)}\cdot t),$ (19)
$\phi_{4}^{(\ell)}(x,t)$ $=\text{coeff}_{4}^{(\ell)}\cdot\cos(\omega_{x}^{(\ell)}\cdot x)\cdot\cos(\lambda_{t}^{(\ell)}\cdot t).$ (20)
Here, $\omega_{x}^{(\ell)}$ and $\lambda_{t}^{(\ell)}$ are trainable frequencies, and $\text{coeff}_{j}^{(\ell)}$ are trainable coefficients for $j=1,2,3,4$ and $\ell=1,\ldots,N$ where $N$ is the number of Fourier feature sets.
@@
  • Heat equation: Initial condition $u(x,y,0)=\sin(20\pi x)\sin(\pi y)$ and homogeneous boundary condition $u(x,y,t)=0$ on $\partial\Omega$ for the unit square domain $0,1]^{2}$ suggest features that satisfy the boundary conditions while capturing the initial structure: $\{\sin(20\pi x),\sin(\pi y)\}$ and their products $\{\sin(20\pi x)\sin(\pi y)\}$
  • Burgers equation: Initial condition $u(x,0)=-\sin(\pi x)$ with homogeneous Dirichlet boundaries u(1,t)=u(1,t)=0u(-1,t)=u(1,t)=0 on domain [1,1][-1,1] provides the feature $\{-\sin(\pi x)\}$.
@@
  • Allen-Cahn equation: Initial condition $u(0,x)=x^{2}\cos(\pi x)$ with periodic boundary conditions u(t,x1)=u(t,x+1)u(t,x-1)=u(t,x+1) on domain [1,1][-1,1] suggests features based on the polynomial-trigonometric structure: $\{x^{2},\cos(\pi x),x^{2}\cos(\pi x)\}$.
  • Navier-Stokes equation: For the back-step flow geometry $0,4]\times[0,2]\setminus([0,2]\times[1,2])$ with no-slip boundary conditions and inlet condition $u_{x}=4y(1-y)$, features are derived from the inlet profile and domain constraints: $\{y(1-y),y,1-y\}$ and polynomial combinations that respect the geometric constraints.
@@ A.1.4 Parameter Initialization
  • Fourier Frequencies: Initialized as $\omega_{x}^{(\ell)}=\lambda_{t}^{(\ell)}=\ell\pi/k$ for $\ell=1,2,\ldots,N$, where $N$ is the number of Fourier feature sets. For homogeneous Dirichlet BCs on $[0,L]$, $k=L$ and for non-Dirichlet BCs, $k$ is set to be simply the identity. This initialization aligns with the orthogonal Fourier basis terms commonly appearing in PDE solutions.
  • Fourier Coefficients: All coefficients $\text{coeff}_{j}^{(\ell)}$ are initialized to 1, providing equal weight to all Fourier modes initially.
@@ A.2.1 PINN and FLS-PINN
PINN: We implement the standard Physics-Informed Neural Network as proposed by Raissi et al. raissi2019physics. This serves as our primary baseline, utilizing the vanilla MLP architecture with coordinate inputs directly fed to the network without any feature engineering or positional encoding. The method relies solely on the network’s inherent ability to learn complex mappings between spatial-temporal coordinates and the solution field through the physics-informed loss function. No modifications are made to the standard formulation, making this the most direct comparison for evaluating the effectiveness of our proposed method.
FLS-PINN: The First Layer Sine PINN represents a modification inspired by the SIREN architecture sitzmann2020implicit, where we replace the $\tanh$ activation function in the first hidden layer with the $\sin$ activation function, while maintaining $\tanh$ activations in all subsequent layers. This hybrid approach aims to leverage the beneficial properties of sinusoidal activations for coordinate-based inputs, particularly their ability to represent high-frequency components, while preserving the stability characteristics of $\tanh$ activations in deeper layers. The first layer weights are initialized using a uniform distribution $\mathcal{U}(-1/d_{\text{in}},1/d_{\text{in}})$ where $d_{\text{in}}$ is the input dimension, following the SIREN initialization scheme, while subsequent layers maintain Xavier initialization. This configuration has been shown to improve convergence behavior for certain classes of PDEs, particularly those involving high-frequency solutions or sharp gradients.
@@ A.2.2 W-PINN
From wang2022and, W-PINN (Weighted Physics-Informed Neural Network) is a variant of PINN that addresses spectral bias and imbalanced convergence rates in multi-term loss functions through adaptive weight calibration. For a PDE defined as:
@@
We refer the reader to wang2021eigenvector for details of these calculations and definitions. The key insight of W-PINN is that the eigenvalues of the NTK matrices $\mathbf{K}_{uu}$ and $\mathbf{K}_{rr}$ characterize the convergence rates of the boundary and residual losses respectively. The method proposes adapting the weights according to:
@@
Implementation Details. In our experiments, we run W-PINN following the setup from wang2022and. The method down-weights the residual loss term to reduce its dominance and improve the fitting of boundary conditions. Specifically, we tested different weighting configurations for the residual loss term from the set $\{10^{-1},10^{-3},10^{-4}\}$ while maintaining the boundary condition weight at the standard value. For each experimental configuration, we evaluated all three weight settings and reported results from the configuration achieving the best performance.
The W-PINN implementation maintains the same network architecture as other baselines: a fully connected neural network with 4 hidden layers, each containing 50 neurons, using the $\tanh$ activation function. Network parameters are initialized using Xavier initialization glorot2010understanding. The adaptive weight computation is performed at regular intervals during training to update the loss function weights based on the current Neural Tangent Kernel eigenvalue analysis.
@@ A.2.3 RBA-PINN
For the RBA-PINN baseline anagnostopoulos2023residualbasedattentionconnectioninformation, we implement the residual-based attention scheme coupled with Fourier feature embeddings (RBA+Fourier configuration) to leverage the method’s demonstrated strengths while maintaining computational efficiency. Based on the original paper’s ablation studies, this configuration represents the most effective combination, where RBA alone achieves $3.16\times 10^{-3}$ relative $L^{2}$ error on the Allen-Cahn equation, while the RBA+Fourier variant improves performance by two orders of magnitude.
Network Architecture. Following the original implementation anagnostopoulos2023residualbasedattentionconnectioninformation, we employ a fully connected neural network with 6 hidden layers and 50 neurons per layer, utilizing the $\tanh$ activation function. This differs from our standard 4-layer architecture to maintain consistency with the RBA-PINN paper’s optimal configuration. Network weights are initialized using Xavier initialization glorot2010understanding.
@@
Fourier Feature Embeddings. To enforce periodic boundary conditions as hard constraints, we implement one-dimensional Fourier feature embeddings following anagnostopoulos2023residualbasedattentionconnectioninformation. For problems with periodic boundary conditions in the spatial domain, the input $x$ is augmented with Fourier features:
@@
where $\langle\cdot\rangle$ denotes the mean operator. The total loss function becomes:
@@ A.2.4 RFF
The Random Fourier Feature PINN (RFF) builds upon the theoretical foundations of Bochner’s theorem and Neural Tangent Kernel (NTK) theory by employing random Fourier feature mappings as coordinate embeddings before the input layer of the neural network tancik2020fourier; wang2021eigenvector. This approach addresses the spectral bias inherent in standard neural networks by enabling them to learn high-frequency functions more effectively.
@@
Implementation Details. Following the methodology established in wang2021eigenvector, we implement RFF using 128 features ($m=64$, resulting in $2m=128$ total features after cosine and sine transformations). The frequency scaling parameter $\sigma$ is chosen based on the expected frequency characteristics of the solution domain. Specifically, we adopt a coordinate-dependent scaling strategy:
@@
The random matrix $B$ is sampled once during initialization and remains fixed throughout training, ensuring deterministic feature mappings. We utilize the random-fourier-features-pytorch library implementation long2021rffpytorch for computational efficiency.
@@
Hyperparameter Selection. The frequency parameter $\sigma$ requires careful tuning based on problem characteristics. Following the analysis in wang2021eigenvector, we conducted preliminary experiments with $\sigma\in\{1,10,50,100,200\}$ for spatial coordinates and $\sigma\in\{1,5,10,20\}$ for temporal coordinates. The selected values ($\sigma=200$ for space, $\sigma=10$ for time) consistently provided the best performance across our benchmark problems.
@@ A.2.5 RBF and RBF-P
RBF (Radial Basis Functions PINN): We implement the RBF-based feature mapping method from Zeng et al. zeng2024rbfpinnnonfourierpositionalembedding, which introduces non-Fourier positional embedding as an alternative to commonly used Fourier-based feature mappings. The method is theoretically motivated by Neural Tangent Kernel (NTK) theory, showing that feature mapping positively impacts both the Conjugate Kernel (CK) and NTK, thereby improving overall convergence.
@@
$\Phi(x)=\frac{\sum_{i=1}^{m}w_{i}\,\phi(\|x-c_{i}\|)}{\sum_{j=1}^{m}\phi(\|x-c_{j}\|)}.$ (36)
where $x\in\mathbb{R}^{n}$ is the input coordinate, $c_{i}\in\mathbb{R}^{m}$ are the RBF centers, and $w_{i}$ are the trainable weights. We employ Gaussian RBFs as the radial basis function
$\phi(r)=e^{-\frac{r^{2}}{\sigma^{2}}},$ (37)
where $r=\|x-c\|$ represents the Euclidean distance between input and RBF center, and $\sigma$ controls the bandwidth.
@@
  • Normalization: The feature mapping includes normalization to ensure $\int K_{\Phi}(x-x^{\prime})dx=1$ and satisfies the symmetry condition for first-order accuracy of the composed kernel approximation.
@@ A.3 Optimizers
As discussed in the main text, we employ a systematic approach to optimizer selection, motivated by recent loss-landscape studies showing that first-order methods like Adam excel at escaping saddle points but struggle in ill-conditioned valleys, while quasi-Newton methods like L-BFGS provide superior local conditioning but are prone to stalling at saddles rathore2024challenges. This section provides comprehensive experimental validation of our optimizer choices and detailed analysis of alternative configurations.
@@ A.3.1 Experimental Configurations
We conducted extensive experiments across multiple optimizer combinations to validate our scheduling choices. All experiments with hybrid Adam and L-BFGS combination optimizers maintain a total training budget of 40k iterations across the tested PDE benchmarks. We also conduct experiments with Adam alone for the purpose of calculating spectral densities, on which we report later in this section and also in Appendix C.2 and C.4. We list the configurations already explained in the main text below and move on to studying other possible configurations and their results.
@@ A.3.2 Experimental Results
Figure 12: Precentage of improvement or deterioration (in L2RE) for each method and each tested PDE when switching from Optimization Schedule (1) to Optimization Schedule (2). It is observed that for SAFE-NET, there is consistent improvement. However, $(\text{Adam}+\text{L-BFGS})^{2}$ doesn’t seem to have any consistent effects on the performances of the rest of the top performing baseline methods as the result varies based on the problem.